OpenStudy (hibs):

(4^13 - 4^12) / (4^11)

OpenStudy (raden):

split into 2 parts, it would becomes 4^13/4^11 - 4^12/4^11 now use the property exponent : a^m/a^n = a^(m-n) continue it ...

OpenStudy (hibs):

So it should be 4^2 - 4^1 =16 - 4 =12 Thank you!

OpenStudy (phi):

another way is factor out 4^12 from the parens (4^13 - 4^12) / (4^11) (4 - 1) 4^12/4^11 3 * 4 12

OpenStudy (hibs):

Thank you, both!